OpenStudy (anonymous):

Sorry if this doesn't make much sense, I'm not familiar with the english mathematical terms. I have a quadratic formula that goes 2x^2+3x+c=0 And I also know that one of the solutions is 1. Now I need to figure out how to determine what 'c' is in this equation. If it helps any, I know that the answer is that 'c' is -5, but would of course like to know how to figure that out.

OpenStudy (amistre64):

you know that 1 is a solution; therefore (x-1) has to be a factor

OpenStudy (amistre64):

(x-1)(kx+n) = 2x^2+3x+c would be a good way to compare

OpenStudy (amistre64):

let me edit that alittle ... dont need thos defulters :) (x-1)(kx+n) = 2x^2+3x+c kx^2 -kx +nx -n = 2x^2+3x+c ; k = 2 by default 2x^2 -2x +nx -n = 2x^2+3x+c 2x^2 +(n-2)x -n = 2x^2+3x+c n-2 = 3

OpenStudy (anonymous):

I'm of course not going to question whether that's right or not, but I'm not entirely sure how that answers the question? As stated before I know that the answer is that 'c = -5', and I don't really understand how you get that from your answer.

OpenStudy (amistre64):

the only way 2 polynomials are equal is if they have the same parts ....

OpenStudy (amistre64):

if 1 is a solution to the setup; then we know that (x-1) is a factor of the setp (x-1), times something; is equal to 2x^2+3x+c let that unknown something be a general linear factor .. (kx+n) (x-1)(kx+n) has to have the same parts as 2x^2+3x+c its just a matter of expanding it out and sorting thru the parts

OpenStudy (amistre64):

(x-1)(kx+n) = x(kx+n) -1(kx+n) = kx^2 +nx -kx -n = kx^2 + (n-k)x -n =2x^2 + 3 x +c comparing parts we have: k=2 n-k = 3 c = -n

OpenStudy (anonymous):

Ahhh, I understand now. Thank you!

OpenStudy (amistre64):

youre welcome :)